# Exploring sub-GeV dark matter via $s$-wave, $p$-wave, and resonance annihilation with CMB data

### v1

@@ IV Numerical results
Table 1: Comparison of constraint power between FIRAS $\mu$-distortion and Planck CMB anisotropies. Note that the $y$-distortion constraint is weaker than those from $\mu$-distortion or CMB anisotropies.
We compare the 95% upper limits on $\mu$-distortion and $y$-distortion from FIRAS with those from CMB anisotropies in Planck. We find that $\mu$-distortion provides a more stringent constraint only for $p$-wave annihilation, due to its larger cross-section at higher redshifts ($5\times 10^{4}\lesssim z\lesssim 2\times 10^{6}$). Therefore, we omit the $y$-distortion analysis. For the resonance scenario, we set the resonance peaks for $\mu$-distortion at $z_{\text{peak}}=10^{6}$ and $y$-distortion at $z_{\text{peak}}=10^{4}$, with DM parameters $m_{\chi}=1~{}\text{GeV}$, $T_{\text{kd}}=m_{\chi}/3$, and $B_{i}B_{f}=0.25$ (which yields a larger cross-section for comparison). However, we find that the resonance effect on the distortion is negligible, as the resonance redshift range is much narrower than that of the distortions, resulting in minimal energy injection. The most stringent limits for the three annihilation scenarios are summarized in Tab. 1.
@@
The left panel of Fig. 2 shows that the PL result is stronger than the MP result, as expected. The MP method averages over all parameters, smoothing out nuisance effects and leading to less stringent constraints. In contrast, the PL method maximizes the likelihood for each parameter, providing sharper and more stringent limits on $\langle\sigma v_{\text{rel}}\rangle$. In the right panel of Fig. 2, a lower mass limit of 150 MeV is set for the $\pi\pi$ final state, based on the rest mass of the $\pi^{\pm}$ (139.57 MeV) and $\pi^{0}$ (134.98 MeV). Our $e^{+}e^{-}$ limits agree with previous studies Cang et al. (2020), while future experiments improve the constraints by approximately a factor of around 3. Note that blackbody spectral distortions limits are not as powerful as the Planck+BAO likelihood for probing the $s$-wave annihilation, even with the most advanced future blackbody spectral distortions missions Chluba (2013a); Fu et al. (2021).
@@
Figure 3: The upper limits projected on the $T_{\rm kd}$ versus $b(1-f_{\nu})$) plane (left) and the $T_{\rm kd}$ versus $b(1-f_{\nu})/m_{\chi}^{2}$ plane (right). The black region represents the last scattering surface, assuming kinetic decoupling occurs after chemical decoupling but before the CMB. The solid lines show the 95% upper limit of FIRAS $\mu$-distortion constraints for $m_{\chi}=10\,\mathrm{MeV}$ (green line) and $m_{\chi}=500\,\mathrm{MeV}$ (magenta line). Their corresponding dashed lines represent the 95% upper limits of $b(1-f_{\nu})$ derived from the Planck+BAO likelihood, by using the MP method. In the right panel, the red solid, black dashed, and blue dash-dotted lines represent the FIRAS, PRISTINE, and PIXIE the 95% upper limit of $\mu$-distortion constraints, respectively.
@@
Fig. 3 shows the constraints for $p$-wave DM annihilation. The left panel displays the relation between $T_{\rm kd}$ and $b(1-f_{\nu})$, while the right panel shows $T_{\rm kd}$ versus $b(1-f_{\nu})/m_{\chi}^{2}$. Here, $f_{\nu}$ is the fraction of DM annihilation energy that goes into neutrinos. For DM annihilation into $e^{+}e^{-}$, $f_{\nu}$ is zero, while for charged pions, $f_{\nu}$ approximates to $0.47$. The black region represents the last scattering surface, assuming kinetic decoupling occurs after chemical decoupling but before the CMB epoch. In the left panel, the FIRAS $\mu$-distortion constraints (solid lines) are stronger than those from the Planck+BAO likelihood (dashed lines), as the $p$-wave cross-section increases with redshift, making the $\mu$-distortion constraint stronger than the $y$-distortion one Li (2024). Similarly, BBN and light-element abundances can yield stronger limits than the FIRAS constraints. However, future missions like PIXIE and Super-PIXIE will surpass BBN limits Braat and Hufnagel (2024). The green and magenta lines show the 95% upper limits for $e^{+}e^{-}$ ($m_{\chi}=10\,\mathrm{MeV}$) and $\pi\pi$ ($m_{\chi}=500\,\mathrm{MeV}$), respectively. Both solid and dashed lines are derived by the MP method. Although the PL result is stronger than the MP result, it remains weaker than the FIRAS constraints.
@@
Fig. 4 shows the 95% upper limits on $B_{i}B_{f}$ for the resonance annihilation scenario, derived from the Planck+BAO likelihood. For a conservative approach, we fix the resonance peak at $z_{\text{peak}}=1100$, which yields the weakest limits compared to those at the range $1100600\leq z_{\text{peak}}<$ 1100, as shown in the right panel of Fig. 1. The limits are obtained using the PL method (solid lines) and the MP method (dashed lines). The left and right panels show results for $e^{+}e^{-}$ and $\pi\pi$ final states, respectively. The values of $x_{\text{kd}}\equiv m_{\chi}/T_{\text{kd}}$ are set to 20 (blue lines), $10^{2}$ (red lines), $10^{3}$ (green lines), and $10^{4}$ (purple lines). The thin cyan dashed lines are derived by future LiteBird and CMB-S4 sensitivities Fu et al. (2021) using the configurations $x_{\text{kd}}=20$ and the MP method. Note that smaller values of $x_{\text{kd}}$ correspond to earlier kinetic decoupling (larger $T_{\text{kd}}$) for the same DM mass. As shown in Eq. (7), larger $T_{\text{kd}}$ results in smaller $\xi$, which strengthens the constraints on $B_{i}B_{f}$. For reference, we compute $\langle\sigma v_{\text{rel}}\rangle$ at $11001+z=z_{\text{peak}}=$ 1100 for a given $B_{i}B_{f}$. Taking $m_{\chi}=1\,\text{GeV}$ and $x_{\text{kd}}=20$ (blue dashed line), the upper limits $B_{i}B_{f}=9\times 10^{-15}$ (${}^{+}e^{-}$) and $B_{i}B_{f}=2\times 10^{-14}$ ($\mu^{+}\mu^{-}$) correspond to annihilation cross-sections $\langle\sigma v_{\text{rel}}\rangle$ of $7.03\times 10^{-28}\,\text{cm}^{3}\,\text{s}^{-1}$ and $1.77\times 10^{-27}\,\text{cm}^{3}\,\text{s}^{-1}$, respectively, at $11001+z=z_{\text{peak}}=$ 1100.
@@ V Conclusion
In this work, we have systematically investigated the constraints imposed by the Planck, BAO and FIRAS likelihood on sub-GeV DM annihilation, considering both $e^{+}e^{-}$ and $\pi\pi$ final states through $s$-wave, $p$-wave, and resonance processes. The comparison of constraint strength between FIRAS $\mu$-distortion and Planck CMB anisotropies is summarized in Table 1, where we note that the $y$-distortion constraint is weaker than those from $\mu$-distortion or CMB anisotropies.
For the $s$-wave annihilation scenario, we have provided constraints for both the $e^{+}e^{-}$ and $\pi\pi$ channels. We find that the PL result is stronger than the MP result, as expected. We also further extend the analysis to include the pion channel and future experiments, such as LiteBird and CMB-S4, both derived using the MP method. Our results for the $e^{+}e^{-}$ channel agree with previous studies, while the inclusion of the $\pi\pi$ channel provides new insights into the constraints on DM annihilation with $m_{\chi}\sim\mathcal{O}(100)\,\mathrm{MeV}$.
@@
For the resonance annihilation scenario, we focused on the case where the resonance peak occurs at $z_{\rm peak}=1100$, which yields the weakest limits compared to the case of $1100600\leq z_{\text{peak}}<$ 1100. Under this assumption, we calculated the constraints on the coupling coefficient $B_{i}B_{f}$ for both $e^{+}e^{-}$ and $\pi\pi$ final states. Our results indicate that the constraints on $\sin\theta$ are significantly more stringent than those from the future DarkSide-50 experiment, highlighting the power of CMB observations in probing DM properties.
In summary, our analysis demonstrates the complementary strengths of $\mu$-distortion and CMB anisotropies constraints in probing sub-GeV DM annihilation. The inclusion of the $\pi\pi$ channel and the use of the PL method provide new and more stringent limits on $s$-wave annihilation. The model-independent inequality for $p$-wave annihilation and the future prospects of PIXIE and Super-PIXIE offer promising avenues. Finally, the resonance scenario, even under the weakest assumptions, provides constraints that surpass those from direct detection experiments.
@@ Appendix A The root-mean-square velocity and $p$-wave annihilation
After DM kinetic decoupling and becoming non-relativistic, its temperature $T_{\chi}\propto(1+z)^{2}$. Using equipartition of energy for ideal gas $\langle v_{\chi}^{2}\rangle=3T_{\chi}/m_{\chi}$, we have
@@
where $z_{\texttt{100}}$ is the redshift when $\langle v_{\chi}^{2}\rangle=v_{\texttt{100}}^{2}$ and $T_{\texttt{100}}$ is the DM temperature at this redshift. By using Eq. (16) and equiparition of energy, we make the substitution
@@ Appendix B The velocity-averaged cross-section near resonance
$\langle\sigma v_{\text{rel}}\rangle=\frac{16\pi^{2}}{(2\pi v^{2}_{\text{rms}}/3)^{3}}\frac{32\pi}{m_{\phi}^{2}\bar{\beta_{i}}}B_{i}B_{f}\int\text{d}\vec{v}_{\text{cm}}v^{2}_{\text{cm}}e^{-3v^{2}_{\text{cm}}/v^{2}_{\text{rms}}}\int\text{d}\vec{v}_{\text{rel}}v^{2}_{\text{rel}}e^{-3v^{2}_{\text{rel}}/(4v^{2}_{\text{rms}})}\frac{\gamma^{2}}{(-\xi+v_{\text{rel}}^{2}/4)^{2}+\gamma^{2}}.$ (23)

### v2

@@ IV Numerical results
Table 1: Comparison of constraining power between FIRAS $\mu$-distortion and Planck CMB anisotropies. Note that the $y$-distortion constraint is weaker than those from $\mu$-distortion or CMB anisotropies.
We compare the 95% upper limits on $\mu$-distortion and $y$-distortion from FIRAS with those from CMB anisotropies in Planck. We find that $\mu$-distortion provides a more stringent constraint only for $p$-wave annihilation, due to its larger cross-section at higher redshifts ($5\times 10^{4}\lesssim z\lesssim 2\times 10^{6}$). Therefore, we omit the $y$-distortion analysis. For the resonance scenario, we set the resonance peaks for $\mu$-distortion at $z_{\text{peak}}=10^{6}$ and $y$-distortion at $z_{\text{peak}}=10^{4}$. However, we find that the resonance effect on the distortion is negligible, as the resonance redshift range is much narrower than that of the distortions, resulting in minimal energy injection. The most stringent limits for the three annihilation scenarios are summarized in Table. 1.
@@
The left panel of Fig. 2 shows that the PL result is stronger than the MP result, as expected. The MP method averages over all parameters, smoothing out nuisance effects and leading to less stringent constraints. In contrast, the PL method maximizes the likelihood for each parameter, providing sharper and more stringent limits on $\langle\sigma v_{\text{rel}}\rangle$. In the right panel of Fig. 2, a lower mass limit of 150 MeV is set for the $\pi\pi$ final state, based on the rest mass of the $\pi^{\pm}$ (139.57 MeV) and $\pi^{0}$ (134.98 MeV). Our $e^{+}e^{-}$ limits agree with previous studies Cang et al. (2020), while future experiments improve the constraints by a factor of approximately 3. Note that blackbody spectral distortion limits are not as powerful as the Planck+BAO likelihood for probing the $s$-wave annihilation, even with the most advanced future blackbody spectral distortions missions Chluba (2013a); Fu et al. (2021).
@@
Figure 3: The upper limits projected on the $T_{\rm kd}$ versus $b(1-f_{\nu})$ plane (left) and the $T_{\rm kd}$ versus $b(1-f_{\nu})/m_{\chi}^{2}$ plane (right). The black region represents the last scattering surface, assuming kinetic decoupling occurs after chemical decoupling but before the CMB. The solid lines show the 95% upper limit of FIRAS $\mu$-distortion constraints for $m_{\chi}=10\,\mathrm{MeV}$ (green line) and $m_{\chi}=500\,\mathrm{MeV}$ (magenta line). Their corresponding dashed lines represent the 95% upper limits of $b(1-f_{\nu})$ derived from the Planck+BAO likelihood, by using the MP method. In the right panel, the red solid, black dashed, and blue dash-dotted lines represent the FIRAS, PRISTINE, and PIXIE the 95% upper limit of $\mu$-distortion constraints, respectively.
@@
Fig. 3 shows the constraints for $p$-wave DM annihilation. The left panel displays the relation between $T_{\rm kd}$ and $b(1-f_{\nu})$, while the right panel shows $T_{\rm kd}$ versus $b(1-f_{\nu})/m_{\chi}^{2}$. Here, $f_{\nu}$ is the fraction of DM annihilation energy that goes into neutrinos. For DM annihilation into $e^{+}e^{-}$, $f_{\nu}$ is zero, while for charged pions, $f_{\nu}$ is approximately $0.47$. The black region represents the last scattering surface, assuming kinetic decoupling occurs after chemical decoupling but before the CMB epoch. In the left panel, the FIRAS $\mu$-distortion constraints (solid lines) are stronger than those from the Planck+BAO likelihood (dashed lines), as the $p$-wave cross-section increases with redshift, making the $\mu$-distortion constraint stronger than the $y$-distortion one Li (2024). Similarly, BBN and light-element abundances can yield stronger limits than the FIRAS constraints. However, future missions like PIXIE and Super-PIXIE will surpass BBN limits Braat and Hufnagel (2024). In addition to the homogeneous DM background contribution, structure formation at $z<50$ enhances the squared number density and velocity dispersion of DM, which could significantly improve constraints on $p$-wave annihilation Diamanti et al. (2013). For 1 GeV DM annihilating into electrons with $T_{\text{kd}}=2.02\,\mathrm{MeV}$, the result from Diamanti et al. (2013) derives a 95% upper limit of $b=1.8\times 10^{-16}\,\text{cm}^{3}\,\text{s}^{-1}$, depending on halo model assumption. In comparison, our model-independent constraints from FIRAS and PIXIE yield less stringent limits of $b<7.6\times 10^{-13}\,\text{cm}^{3}\,\text{s}^{-1}$ and $b<9.9\times 10^{-16}\,\text{cm}^{3}\,\text{s}^{-1}$, respectively, without relying on specific DM clustering assumption. The green and magenta lines show the 95% upper limits for $e^{+}e^{-}$ ($m_{\chi}=10\,\mathrm{MeV}$) and $\pi\pi$ ($m_{\chi}=500\,\mathrm{MeV}$), respectively. Both solid and dashed lines are derived by the MP method. Although the PL result is stronger than the MP result, it remains weaker than the FIRAS constraints.
@@
Fig. 4 shows the 95% upper limits on $B_{i}B_{f}$ for the resonance annihilation scenario, derived from the Planck+BAO likelihood. For a conservative approach, we fix the resonance peak at $z_{\text{peak}}=1100$, which yields the weakest limits compared to those at the range $1100600\leq z_{\text{peak}}<$ 1100, as shown in the right panel of Fig. 1. The limits, obtained using the PL method (solid lines) and the MP method (dashed lines), are shown for the $e^{+}e^{-}$ (left panel) and $\pi\pi$ (right panel) final states. The values of $x_{\text{kd}}\equiv m_{\chi}/T_{\text{kd}}$ are set to 20 (blue lines), $10^{2}$ (red lines), $10^{3}$ (green lines), and $10^{4}$ (purple lines). The thin cyan dashed lines are derived by future LiteBIRD and CMB-S4 sensitivities Fu et al. (2021) using the configurations $x_{\text{kd}}=20$ and the MP method. Note that smaller values of $x_{\text{kd}}$ correspond to earlier kinetic decoupling (larger $T_{\text{kd}}$) for the same DM mass. As shown in Eq. (7), larger $T_{\text{kd}}$ results in smaller $\xi$, which strengthens the constraints on $B_{i}B_{f}$. For reference, we compute $\langle\sigma v_{\text{rel}}\rangle$ at $11001+z=z_{\text{peak}}=$ 1100 for a given $B_{i}B_{f}$. Taking $m_{\chi}=1\,\text{GeV}$ and $x_{\text{kd}}=20$ (blue dashed line), the upper limits $B_{i}B_{f}=9\times 10^{-15}$ (${}^{+}e^{-}$) and $B_{i}B_{f}=2\times 10^{-14}$ ($\mu^{+}\mu^{-}$) correspond to annihilation cross-sections $\langle\sigma v_{\text{rel}}\rangle$ of $7.03\times 10^{-28}\,\text{cm}^{3}\,\text{s}^{-1}$ and $1.77\times 10^{-27}\,\text{cm}^{3}\,\text{s}^{-1}$, respectively, at $11001+z=z_{\text{peak}}=$ 1100.
@@ V Conclusion
In this work, we have systematically investigated the constraints imposed by the Planck, BAO, and FIRAS likelihood on sub-GeV DM annihilation, considering both $e^{+}e^{-}$ and $\pi\pi$ final states through $s$-wave, $p$-wave, and resonance processes. The comparison of constraint strength between FIRAS $\mu$-distortion and Planck CMB anisotropies is summarized in Table 1, where we note that the $y$-distortion constraint is weaker than those from $\mu$-distortion or CMB anisotropies.
For the $s$-wave annihilation scenario, we have provided constraints for both the $e^{+}e^{-}$ and $\pi\pi$ channels. We find that the PL result is stronger than the MP result, as expected. We have further extended the analysis to include the pion channel and future experiments, such as LiteBIRD and CMB-S4, both derived using the MP method. Our results for the $e^{+}e^{-}$ channel agree with previous studies, while the inclusion of the $\pi\pi$ channel provides new insights into the constraints on DM annihilation with $m_{\chi}\sim\mathcal{O}(100)\,\mathrm{MeV}$.
@@
For the resonance annihilation scenario, we focused on the case where the resonance peak occurs at $z_{\rm peak}=1100$, which yields the weakest limits compared to the case of $1100600\leq z_{\text{peak}}<$ 1100. Under this assumption, we calculated the constraints on the coupling coefficient $B_{i}B_{f}$ for both $e^{+}e^{-}$ and $\pi\pi$ final states. Our results indicate that the constraints on $\sin\theta$ are significantly more stringent than those from the DarkSide-50 experiment, highlighting the power of CMB observations in probing DM properties.
In summary, our analysis demonstrates the complementary strengths of $\mu$-distortion and CMB anisotropies constraints in probing sub-GeV DM annihilation. The inclusion of the $\pi\pi$ channel and the use of the PL method provides new and more stringent limits on $s$-wave annihilation. The model-independent inequality for $p$-wave annihilation and the future prospects of PIXIE and Super-PIXIE offer promising avenues. Finally, the resonance scenario, even under the weakest assumptions, provides constraints that surpass those from direct detection experiments.
@@ Appendix A The root-mean-square velocity and $p$-wave annihilation
After DM kinetic decoupling and becoming non-relativistic, its temperature $T_{\chi}\propto(1+z)^{2}$. Using equipartition of energy for an ideal gas, $\langle v_{\chi}^{2}\rangle=3T_{\chi}/m_{\chi}$, we have
@@
where $z_{\texttt{100}}$ is the redshift when $\langle v_{\chi}^{2}\rangle=v_{\texttt{100}}^{2}$ and $T_{\texttt{100}}$ is the DM temperature at this redshift. By using Eq. (16) and equipartition of energy, we make the substitution
@@ Appendix B The velocity-averaged cross-section near resonance
$\langle\sigma v_{\text{rel}}\rangle=$ $\frac{16\pi^{2}}{(2\pi v^{2}_{\text{rms}}/3)^{3}}\frac{32\pi}{m_{\phi}^{2}\bar{\beta_{i}}}B_{i}B_{f}\int\text{d}\vec{v}_{\text{cm}}v^{2}_{\text{cm}}e^{-3v^{2}_{\text{cm}}/v^{2}_{\text{rms}}}$ (23)
$\int\text{d}\vec{v}_{\text{rel}}v^{2}_{\text{rel}}e^{-3v^{2}_{\text{rel}}/(4v^{2}_{\text{rms}})}\frac{\gamma^{2}}{(-\xi+v_{\text{rel}}^{2}/4)^{2}+\gamma^{2}}.$